OpenStudy (anonymous):

Need help with this problem that needs to be solved using substitution: 3/4x-1/12y=-5/2 3/8x-1/6y=-1/2 I came up with several different answers and really don't know what to do...I tried multiplying it with 12 or 24 and it just doesn't makes sense..please help, thanks.....

OpenStudy (amistre64):

multiply eq1 by 12 ; and eq 2 by 24 to clear the fractions

OpenStudy (anonymous):

Multiply Both sides of the first equation with 12xy and both sides of second equation with 24xy. The new equations are 9y - x = -30xy 9y - 4x = -12xy Now solve for x and y. Ans. y= -1/6, x = -1/4

OpenStudy (amistre64):

thats also plausible; assuming x and y are denominators :)

OpenStudy (amistre64):

usually this is a typo and means x and y are simply multiple of the fractions.

OpenStudy (anonymous):

Thank you for your help - I came up with some weird number though..thanks again :)

OpenStudy (anonymous):

:)